OpenStudy (anonymous):

Please Help Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1) y=1/3x - 3 y=1/3x + 13 y = −3x − 3 y = −3x + 17

OpenStudy (shamim):

I think its s question of math

OpenStudy (anonymous):

Yes

OpenStudy (shamim):

3x+y=7 y=-3x+7

OpenStudy (anonymous):

you mean 17 instead of 7?

OpenStudy (shamim):

So slope of ur given line is m=-3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ty

OpenStudy (shamim):

U know m1*m2=-1 -3*m2=-1 m2=1/3

OpenStudy (shamim):

So the slope of ur required line is 1/3

OpenStudy (anonymous):

So its going to be the first one?

OpenStudy (shamim):

Ya

OpenStudy (anonymous):

ty

OpenStudy (shamim):

So the equation of ur required line will b y-y1=m(x-x1)

OpenStudy (anonymous):

Yes im sure of that

OpenStudy (shamim):

Here m=1/3 x1=6 y1-1

OpenStudy (shamim):

So y-(-1)=1/3(x-6)

OpenStudy (shamim):

Now finish it

OpenStudy (anonymous):

the last one

OpenStudy (shamim):

No

OpenStudy (shamim):

y+1=1/3x-2 y=1/3x-3

OpenStudy (shamim):

Did u get it